OpenStudy (anonymous):

If f(x)=3/x+2-radical x-3 complete the following statement. The domain for f(x) is all real number ___than or equal to 3

OpenStudy (freckles):

is this \[f(x)=\frac{3}{x}+2-\sqrt{x}-3 \\ \text{ or } f(x)=\frac{3}{x+2}-\sqrt{x}-3 \\ \text{ or } f(x)=\frac{3}{x}+2-\sqrt{x-3} \\ \text{ or } f(x)=\frac{3}{x+2}-\sqrt{x-3}\]

OpenStudy (freckles):

think the following for any of the above: you want the inside of the square root to be positive or 0 but you do not want to divide by 0